OpenStudy (anonymous):

Factorize 8-12x+6x^2-x^3 ...?

OpenStudy (anonymous):

12^x5 duh

jhonyy9 (jhonyy9):

8-x3=(2-x)(4+2x+x2) 6x2-12x=-6x(2-x) so hence your exercise is equal (2-x)(4+2x+x2)-6x(2-x)=(2-x)(4+2x+x2-6x)= (2-x)(4-4x+x2)=(2-x)(2-x)2 =(2-x)3 (2-x)2 is (2-x) on exponent 2 (2-x)3 is (2-x) on exponent 3